### 23 Jan 2024 v2

# Spatial-Temporal Large Language Model for Traffic Prediction

Abstract
Traffic prediction, a critical component for intelligent transportation systems, endeavors to foresee future traffic at specific locations using historical data. Although existing traffic prediction models often emphasize developing complex neural network structures, their accuracy has not seen improvements accordingly. Recently, Large Language Models (LLMs) have shown outstanding capabilities in time series analysis. Differing from existing models, LLMs progress mainly through parameter expansion and extensive pre-training while maintaining their fundamental structures. In this paper, we propose a Spatial-Temporal Large Language Model (ST-LLM) for traffic prediction. Specifically, ST-LLM redefines the timesteps at each location as tokens and incorporates a spatial-temporal embedding module to learn the spatial location and global temporal representations of tokens. Then these representations are fused to provide each token with unified spatial and temporal information. Furthermore, we propose a novel partially frozen attention strategy of the LLM, which is designed to capture spatial-temporal dependencies for traffic prediction. Comprehensive experiments on real traffic datasets offer evidence that ST-LLM outperforms state-of-the-art models. Notably, the ST-LLM also exhibits robust performance in both few-shot and zero-shot prediction scenarios.
1 Introduction
Traffic prediction, which aims to predict future traffic features like flow at specific locations using historical data, is a crucial component for intelligent transportation systems Li et al. (2023); Jiang et al. (2023); Chang et al. (2023); Li et al. (2023); Gong et al. (2023). This prediction is instrumental in optimizing traffic management Miao et al. (2024); Zhou et al. (2024) and scheduling public transportation Jin et al. (2023a); Wang et al. (2020). For instance, accurately predicting bike flow benefits the transportation department in optimizing bike management. Similarly, taxi flow forecasting is vital for vehicle companies, as it enables them to efficiently allocate and schedule vehicles to satisfy expected demand.
The evolution of traffic prediction has seen a shift from traditional time series models to deep learning techniques Campos et al. (2023); Chang et al. (2023); Yuan et al. (2018); Kumar and Vanajakshi (2015). Initially, time series models such as the Autoregressive Integrated Moving Average and Kalman Filter were adapted for their fit with time series data. However, these models are not good at capturing the spatial-temporal dependencies within traffic data, leading to deep learning solutions using convolutional neural networks (CNNs) for spatial and recurrent neural networks (RNNs) for temporal dependencies Yin et al. (2021); Shen et al. (2018); Yuan et al. (2018). Despite these advancements, the non-Euclidean spatial structure and the complex periodicity of traffic data present challenges for CNNs and RNNs in capturing spatial and temporal dependencies well.
Graph convolutional network (GCN) based models gained popularity for their ability to model local spatial dependencies Li et al. (2023); Bai et al. (2020, 2020); Wu et al. (2019); Li et al. (2018); Yu et al. (2018). However, these models often encountered over-smoothing issues, which limits their ability to capture global spatial patterns. This shortcoming prompted a shift to attention-based models, which effectively model dynamic spatial correlations without depending on an adjacency matrix Jiang et al. (2023); Guo et al. (2022); Lin et al. (2020). These attention-based approaches have since emerged as a leading trend, offering a superior ability of handling spatial-temporal dependencies in traffic prediction Zheng et al. (2020); Guo et al. (2019). Nevertheless, with this evolution, the structures of existing traffic prediction models have become progressively complex.
Foundation models, including large language models (LLMs), have advancements in fields such as computer vision Ko et al. (2023); Yu et al. (2023) and natural language processing Ramezani and Xu (2023); Maynez et al. (2023). More recently, LLMs also have shown superb performance on time series analysis Cao et al. (2023); Jin et al. (2023b). Compared with the complex designs of existing predictive models, LLMs primarily evolve through the expansion of parameters and pre-training while maintaining their foundational model structure. Existing LLM-based prediction methods tend to focus on the temporal aspect of data in the traffic prediction tasks Zhou et al. (2023); Jin et al. (2023b); Cao et al. (2023) and often overlook the spatial aspect. However, in traffic prediction, the spatial variables are strongly correlated and the spatial dimension also proves to be important Lablack and Shen (2023); Wen et al. (2023). For example, a common problem setting is to use traffic data from the previous twelve timesteps to predict traffic for the next twelve timesteps at hundreds of spatial locations Li et al. (2018) - in this case, more spatial data than temporal data can be leveraged. In our study, we refine the timesteps of a spatial location as a token and model the global temporal dependencies across all these tokens so as to emphasize the spatial aspects.
Moreover, LLMs are notable for their ability to transfer knowledge across domains Rasul et al. (2023); Nate Gruver and Wilson (2023), such as the pretrained transformer (FPT) LLM Zhou et al. (2023). While the FPT LLM is effective in time series analysis tasks, it shows less optimal performance in long-term prediction tasks like traffic prediction. The possible reason is that FPT struggles to bridge the domain gap between language data and traffic data. To fill this gap, we propose a partially frozen attention (PFA) LLM, specifically designed to enhance prediction accuracy in traffic prediction. By partially freezing the multi-head attention, the LLM can adapt to traffic prediction while preserving the foundational knowledge acquired during pre-training.
In summary, we propose a novel framework for traffic prediction termed Spatial-Temporal Large Language Model (ST-LLM). In this model, we define timesteps at a location as a token. These tokens transform a specialized spatial-temporal embedding layer, which is designed to emphasize spatial and temporal correlations such as spatial correlations and temporal patterns. Furthermore, we fuse the spatial-temporal embeddings of each token. This fusion process plays a pivotal role in aggregating spatial and temporal embeddings into a unified representation. Following this, we introduce the partially frozen attention LLM, a novel strategy tailored for LLMs to effectively capture the spatial-temporal dependencies in traffic prediction. Extensive experiments on real-world traffic datasets have validated the efficacy of ST-LLM. The key contributions of this paper are summarized as follows:
We propose a novel Spatial-Temporal Large Language Model (ST-LLM) for traffic prediction, which defines timesteps at a location as a token and embeds each token by a spatial-temporal embedding layer. We fuse the spatial-temporal embeddings of these tokens uniformly and adapt the LLMs for capturing global spatial-temporal dependencies.
A novel strategy within the LLM, named partially frozen attention, is proposed to enhance the model in traffic prediction. By partially freezing the multi-head attention, the ST-LLM is adapted to capture global spatial-temporal dependencies between tokens for different traffic prediction tasks.
Extensive experiments are conducted on real traffic datasets to show the superior performance achieved by our ST-LLM across various settings. Moreover, the few-shot and zero-shot prediction results highlight the ST-LLM’s capability for intra-domain and inter-domain knowledge transfer.
2 Related Work
2.1 Large Language Models for Time Series Analysis
Recently large language models (LLMs) have shown superb performance on time series analysis tasks, such as prediction Cao et al. (2023), classification Jin et al. (2023b), anomaly detection Zhou et al. (2023), imputation Chen et al. (2023), few-shot learning Sun et al. (2023), and zero-shot learning Nate Gruver and Wilson (2023). For instance, TEMPO-GPT combines prompt engineering and seasonal trend decomposition in the generative pre-trained transformer (GPT) Cao et al. (2023). TIME-LLM reprograms an LLM for time series forecasting, and the backbone language model remains intact Jin et al. (2023b). OFA employs a frozen GPT2 model across various key tasks in time series analysis Zhou et al. (2023), the authors conclude that the LLM performs better on tasks of time series analysis. TEST generates various embedding for time series tokens and executes time series forecasting and classification tasks Sun et al. (2023). The above model only models the temporal dimension of the data and ignores the spatial dimension. GATGPT integrates the graph attention network and GPT for spatial-temporal imputation Chen et al. (2023). However, it directly overlooks the temporal representation. As of now, the technique for effectively embedding time series data that encompasses both spatial and temporal representations before inputting it into LLMs is not well-defined.
2.2 Traffic Prediction
Traffic prediction aims to predict future traffic features based on historical traffic data, which is crucial for intelligent transportation systems Ye et al. (2021); Shao et al. (2022); Miao et al. (2023). Traditionally, the traffic prediction model is based on statistics, such as ARIMA and Kalman filter Kumar and Vanajakshi (2015); Chang et al. (2023). However, these models do not perform well due to the inherent spatial-temporal dependencies of traffic data. Later, numerous efforts have been dedicated to advancing traffic prediction techniques through the development of various neural network-based models. First, convolutional neural networks (CNNs) were applied to traffic data to capture spatial dependencies in the data Shen et al. (2018); Yuan et al. (2018). Given that CNNs are primarily designed for regular, grid-like urban areas, they encounter challenges when dealing with the non-Euclidean spatial structure of traffic data. Then, graph convolutional network (GCN) based models became popular due to their permutation-invariance, local connectivity, and compositionally Wu et al. (2019); Bai et al. (2020). More recently, attention-based models have emerged as a dominant trend Guo et al. (2019); Zheng et al. (2020); Guo et al. (2022); Jiang et al. (2023). Without taking the adjacency matrix into account, attention-based models can still model dynamic spatial correlation more effectively than GCN-based models Liu et al. (2023a). However, the structures of these models are becoming increasingly sophisticated.
3 Problem Definition
Definition 3.1 (Traffic Feature). We denote the traffic data as a tensor , where is the number of timesteps, is the number of spatial stations, and is the feature. For example, represents the traffic pick-up or drop-off flow.
Definition 3.2 (Traffic Prediction). Given the historical traffic feature of timesteps , the objective is to learn a function with parameter to predict traffic feature of on the following timesteps . That is,
where .
4 Methodology
In this section, we provide a detailed elaboration of the proposed ST-LLM and its components.
4.1 Overview
The Spatial-Temporal Large Language Model (ST-LLM) framework, as depicted in Figure 1, integrates a spatial-temporal embedding layer, a fusion convolution layer, an LLM layer, and a regression convolution layer. Initially, the historical traffic data is denoted as , which contains tokens of spatial locations. The is processed through the spatial-temporal embedding layer, which extracts the token embedding of historical timestamps, spatial embedding, and temporal embedding, as , , and , respectively. A fusion convolution then integrates these representations into a unified way . Subsequently, the is input into a PFA LLM that encompasses layers, where the multi-head attention and feed forward layers in the first layers are frozen to preserve the pretrained knowledge and the multi-head attention layers in the last layers are unfrozen to enhance the model’s focus on capturing the spatial-temporal dependencies between tokens, resulting in the output . Finally, the regression convolution layer takes and predicts the following traffic data, denoted as .
4.2 Spatial-Temporal Embedding and Fusion
We aim to modify LLMs that have already been trained for traffic prediction tasks. We refine the timestamps at each location of traffic data as tokens. The spatial-temporal embedding layer is designed to transform the tokens into spatial-temporal representations that align with the LLMs. These representations including spatial correlations, hour-of-day, day-of-week patterns, and token information.
We embed the tokens through a pointwise convolution, where the input data is transformed into the embedding :
where represents the token embedding. denotes the pointwise convolution operation using filters that have a kernel size. is the input data, is the hidden dimension. represents the learnable parameters of the pointwise convolution.
To preserve the temporal information in the tokens, we utilize a linear layer to encode the input data into separate embeddings for the hour-of-day and day-of-week temporal embeddings. We perform absolute positional encoding for each traffic data at the “day” and “week” resolutions, and the generated positional encodings are and . The hour-of-day embedding and day-of-week embedding are calculated as follows:
and are the learnable parameter embeddings for the hour-of-day and day-of-week, respectively. By adding these two embeddings, we obtain the temporal representation .
To represent spatial correlations among token pairs, we design an adaptive embedding of tokens, :
where denotes the activation function, and are the learnable parameter embeddings.
Subsequently, we introduce a fusion convolution (FConv) to integrate the token, spatial, and temporal embeddings to represent the spatial-temporal embedding of each token:
where , denotes concatenation, and represents the learnable parameters of the FConv.
4.3 Partially Frozen Attention (PFA) LLM
The frozen pretrained transformer (FPT) has demonstrated effectiveness in a variety of downstream tasks across non-language modalities Lu et al. (2022). However, its performance is less optimal in tasks requiring both short-term and long-term predictions, such as traffic prediction Zhou et al. (2023). To address this limitation, we propose a partially frozen attention (PFA) LLM, specifically designed to enhance prediction accuracy in traffic prediction.
The difference between the FPT and our PFA primarily lies in the configuration of frozen attention layers. In the FPT framework, both the multi-head attention and feed-forward layers are frozen during training, as these layers contain the most significant portion of the learned knowledge within the LLM. In the PFA, we maintain the first layers identical to the FPT, but crucially, we unfreeze the last multi-head attention layers since the attentions offer effective handling of spatial-temporal dependencies in data. Consequently, our ST-LLM can adapt to traffic prediction while preserving the foundational knowledge acquired during pre-training.
Furthermore, our ST-LLM inverts the traditional calculation dimension from temporal to spatial. This inversion is intentional and aligns with the operation of the partially frozen layers. By focusing on spatial dimensions, our model captures global correlations more effectively than if we were to concentrate solely on temporal aspects. This shift is particularly relevant in the context of traffic prediction, where spatial dynamics play a critical role in determining flow patterns.
To elaborate, the PFA LLM generates a probabilistic distribution of the new representations based on the sequence of existing embedded representations . In the first layers of the LLM:
where the range of is from to , and . represent the positional encoding. represents the intermediate representation of the layer after applying the frozen multi-head attention mechanism (MHA) and the first unfrozen layer normalization (LN). symbolizes the final representation of the layer after applying both the unfrozen LN and frozen feed-forward network (FFN).
In the last layers of the LLM, we unfreeze the to adapt the ST-LLM for capturing spatial-temporal dependencies of traffic data:
where represents the intermediate representation of the layer after applying the unfrozen MHA and the second frozen LN. denotes the final output of the layer after applying both the unfrozen LN and frozen FFN, with the MHA being unfrozen.
After the PFA LLM, we design a regression convolution (RConv) to predict the traffic features on the following timesteps:
where , and represents the learnable parameters of the regression convolution.
The loss function of ST-LLM is established as follows:
where is the predicted traffic feature. is the ground truth. represents the L2 regularization term, which helps control overfitting. is a hyperparameter.
Dataset | NYCTaxi Pick-up | NYCTaxi Drop-off | CHBike Pick-up | CHBike Drop-off | ||||||||||||
Metric | MAE | RMSE | MAPE | WAPE | MAE | RMSE | MAPE | WAPE | MAE | RMSE | MAPE | WAPE | MAE | RMSE | MAPE | WAPE |
DCRNN | 5.40 | 9.71 | 35.09% | 20.43% | 5.19 | 9.63 | 37.78% | 19.82% | 2.09 | 3.30 | 54.22% | 42.26% | 1.96 | 2.94 | 51.42% | 39.61% |
STGCN | 5.71 | 10.22 | 36.51% | 21.62% | 5.38 | 9.60 | 39.12% | 20.55% | 2.08 | 3.31 | 53.63% | 42.08% | 2.01 | 3.07 | 50.45% | 40.62% |
ASTGCN | 7.43 | 13.84 | 47.96% | 28.04% | 6.98 | 14.70 | 45.48% | 26.60% | 2.76 | 4.45 | 64.23% | 55.71% | 2.79 | 4.20 | 69.88% | 56.49% |
GWN | 5.43 | 9.39 | 37.79% | 20.55% | 5.03 | 8.78 | 35.63% | 19.21% | 2.04 | 3.20 | 53.08% | 40.95% | 1.95 | 2.98 | 50.30% | 39.43% |
AGCRN | 5.79 | 10.11 | 40.40% | 21.93% | 5.45 | 9.56 | 40.67% | 20.81% | 2.16 | 3.46 | 56.35% | 43.69% | 2.06 | 3.19 | 51.91% | 41.78% |
GMAN | 5.43 | 9.47 | 34.39% | 20.42% | 5.09 | 8.95 | 35.00% | 19.33% | 2.20 | 3.35 | 57.34% | 44.06% | 2.09 | 3.00 | 54.82% | 42.00% |
ASTGNN | 5.90 | 10.71 | 40.15% | 22.32% | 6.28 | 12.00 | 49.78% | 23.97% | 2.37 | 3.67 | 60.08% | 47.81% | 2.24 | 3.35 | 57.21% | 45.27% |
STG-NCDE | 6.24 | 11.25 | 43.20% | 23.46% | 5.38 | 9.74 | 40.45% | 21.37% | 2.15 | 3.97 | 55.49% | 61.38% | 2.28 | 3.42 | 60.96% | 46.06% |
DGCRN | 5.44 | 9.82 | 35.78% | 20.58% | 5.14 | 9.39 | 35.09% | 19.64% | 2.06 | 3.21 | 54.06% | 41.51% | 1.96 | 2.93 | 51.99% | 39.70% |
OFA | 5.82 | 10.42 | 36.67% | 22.00% | 5.60 | 10.14 | 37.39% | 21.36% | 2.06 | 3.21 | 53.55% | 41.70% | 1.96 | 2.97 | 49.64% | 39.68% |
GATGPT | 5.92 | 10.55 | 37.83% | 22.39% | 5.66 | 10.39 | 37.36% | 21.60% | 2.07 | 3.23 | 52.54% | 41.70% | 1.95 | 2.94 | 49.26% | 39.43% |
GCNGPT | 6.58 | 12.23 | 40.19% | 24.88% | 6.64 | 12.24 | 42.46% | 25.32% | 2.37 | 3.80 | 56.24% | 47.66% | 2.24 | 3.48 | 51.05% | 45.37% |
LLAMA2 | 5.35 | 9.48 | 41.32% | 20.27% | 5.66 | 10.74 | 47.47% | 21.63% | 2.10 | 3.37 | 56.63% | 42.49% | 1.99 | 3.03 | 55.23% | 40.28% |
ST-LLM | 5.29 | 9.42 | 33.55% | 20.03% | 5.07 | 9.07 | 33.34% | 19.18% | 1.99 | 3.08 | 53.54% | 40.19% | 1.89 | 2.81 | 49.50% | 38.27% |
5 Experiments
5.1 Datasets
This section details the datasets employed to examine the predictive performance of the ST-LLM and baselines, with real-world traffic data from NYCtaxi Ye et al. (2021), CHBike Jiang et al. (2023).
NYCTaxi. The NYCTaxi dataset comprises over 35 million taxi trips in New York City (NYC), systematically categorized into 266 virtual stations. Spanning three months from April 1st to June 30th, 2016, it includes 4,368 timesteps, each representing a half-hour interval.
CHBike. Consisting of approximately 2.6 million Citi bike orders, the CHBike dataset reflects the usage of the bike-sharing system in the same period as the NYCTaxi dataset, from April 1st to June 30th, 2016. After filtering out stations with few orders, it focuses on the 250 most frequented stations. The dataset aligns with NYCTaxi in terms of time, covering 4,368 timesteps with each timestep representing a 30-minute interval.
LLM | No Pretrain | Full Layer | Full Tuning | FPT | PFA | ||||||||||
Metric | MAE | RMSE | WAPE | MAE | RMSE | WAPE | MAE | RMSE | WAPE | MAE | RMSE | WAPE | MAE | RMSE | WAPE |
CHBike Drop-off | 1.92 | 2.86 | 38.84% | 1.91 | 2.83 | 38.63% | 1.90 | 2.82 | 38.28% | 1.92 | 2.86 | 38.90% | 1.89 | 2.81 | 38.27% |
CHBike Pick-up | 2.03 | 3.14 | 40.87% | 2.02 | 3.12 | 40.62% | 2.01 | 3.11 | 40.43% | 2.07 | 3.25 | 41.65% | 1.99 | 3.08 | 40.19% |
LLM | NYCTaxi Pick-up | NYCTaxi Drop-off | CHBike Pick-up | CHBike Drop-off | ||||||||||||
MAE | RMSE | MAPE | WAPE | MAE | RMSE | MAPE | WAPE | MAE | RMSE | MAPE | WAPE | MAE | RMSE | MAPE | WAPE | |
OFA | 6.49 | 12.12 | 46.74% | 24.54% | 6.27 | 12.10 | 45.23% | 23.92% | 2.20 | 3.59 | 57.52% | 44.40% | 2.06 | 3.17 | 55.96% | 41.63% |
GATGPT | 7.02 | 13.09 | 50.19% | 26.54% | 6.84 | 13.27 | 56.15% | 26.09% | 2.59 | 4.41 | 56.23% | 52.20% | 2.50 | 4.07 | 56.36% | 50.64% |
GCNGPT | 10.31 | 18.82 | 59.41% | 39.02% | 9.25 | 19.50 | 56.77% | 35.28% | 2.73 | 4.44 | 56.93% | 55.20% | 2.79 | 4.65 | 61.85% | 56.28% |
LLAMA2 | 5.81 | 10.16 | 41.82% | 21.99% | 5.59 | 9.90 | 40.58% | 21.35% | 2.24 | 3.58 | 59.47% | 45.20% | 2.11 | 3.23 | 54.44% | 42.75% |
ST-LLM | 5.40 | 9.63 | 33.36% | 20.45% | 5.54 | 9.84 | 39.56% | 21.14% | 2.07 | 3.23 | 55.68% | 41.85% | 1.93 | 2.88 | 52.75% | 39.21% |
5.2 Baselines
We compare our ST-LLM with the following baseline models, which cover different types of methods and have been widely used for spatio-temporal prediction tasks. We compare ST-LLM with the following 10 baselines belonging to two classes. (1) GNN-based models: DCRNN Li et al. (2018), STGCN Yu et al. (2018), GWN Wu et al. (2019), AGCRN Bai et al. (2020), STG-NCDE Choi et al. (2022), DGCRN Li et al. (2023). (2) Attention-based models: ASTGCN Guo et al. (2019), GMAN Zheng et al. (2020), ASTGNN Guo et al. (2022). (3) LLM-based models: OFA Zhou et al. (2023), GATGPT Chen et al. (2023), GCNGPT, and LLAMA2.
5.3 Implementations
Aligning with contemporary practices, we divided the NYCTaxi and CHBike datasets into training, validation, and test sets using a 6:2:2 ratio. We set the historical timesteps and the future timesteps to 12 each, enabling multi-step traffic prediction. is set at 7 to represent a week’s seven days. is 48, with each timestep spanning 30 minutes. The experiments were carried out on a system incorporating NVIDIA A100 GPUs, each with 40GB of memory. For training LLM-based models, we used the Ranger optimizer with a learning rate of 0.001, while GCN and attention-based models employed the Adam optimizer, also set at a 0.001 learning rate. The LLMs used are GPT2 and LLAMA2 7B. We configured GPT2 with six layers Zhou et al. (2023), and LLAMA2 with eight layers Jin et al. (2023b). The experiments were conducted with a batch size of 64 over 200 training epochs. Note that the experimental results are averaged across all prediction timesteps.
5.4 Evaluation Metrics
Four metrics were used for evaluating the models: mean absolute error (MAE), mean absolute percentage error (MAPE), root mean squared error (RMSE), and weighted absolute percentage error (WAPE). MAE and RMSE quantify absolute errors, while MAPE and WAPE assess relative errors. In all metrics, lower values indicate superior prediction performance:
where is the number of predicted values.
5.5 Results and Analysis
The comparison results with baselines are shown in Tables 1. We can make the following observations. (1) LLM-based methods yield superior prediction results, with ST-LLM exhibiting the most effective performance. The ST-LLM outperforms other LLMs in six traffic prediction scenarios, demonstrating its superior accuracy in handling diverse traffic data across various datasets. (2) OFA and LLAMA2 are competent but surpassed by ST-LLM which achieves a 22.5% average MAE improvement over OFA and 20.8% over LLAMA2. This may be due to OFA’s ineffective traffic data embedding, making it difficult for LLMs to understand the spatial-temporal dependencies between data. Despite LLAMA2’s larger size and complexity, it doesn’t directly translate to better traffic prediction than ST-LLM. GATGPT and GCNGPT do not extract temporal representations of traffic data to influence the LLM to capture spatial-temporal dependencies. (3) Attention-based models like ASTGNN and GMAN exhibit varied performance across different datasets. They performed quite well in some cases but were always inferior to ST-LLM. This variability could be attributed to the limitations of traditional attention mechanisms in handling complex spatial-temporal embeddings, especially when compared to large language models. (4) GNN-based Models such as GWN and DGCRN demonstrate competitive performance, particularly in specific metrics, but still cannot outperform ST-LLM. This suggests that while GNNs effectively capture spatial dependencies, their temporal analysis capabilities might not be as advanced as the ST-LLM, which limits their overall performance.
Figure 2 illustrates the trade-off between inference time and MAE for various LLMs on NYCTaxi and CHBike Drop-off datasets. Notably, LLAMA2 is excluded from this comparison due to its substantially longer inference time relative to other LLMs. On the NYCTaxi Drop-off dataset, ST-LLM achieves the lowest MAE while maintaining competitive inference times. Following closely is OFA, whose inference time is similar to ST-LLM, albeit with a slightly higher MAE. This suggests that spatial-temporal embedding and PFA do not slow down the inference speed of the LLM, yet enhance prediction accuracy. GATGPT and GCNGPT exhibit longer inference times and higher MAEs than ST-LLM, with GCNGPT being the slowest. This could be attributed to the fact that while GCN generally has a simpler structure, their combination with GPT might introduce extra computational complexity. This also implies that GAT’s attention mechanism could be more efficient when combined with GPT. Similar trends are observed in the CHBike Drop-off dataset. In summary, ST-LLM stands out as the model providing the best balance between inference speed and predictive accuracy across both datasets.
LLM | OFA | GATGPT | GCNGPT | LLMAM2 | ST-LLM | ||||||||||
MAE | RMSE | MAPE | MAE | RMSE | MAPE | MAE | RMSE | MAPE | MAE | RMSE | MAPE | MAE | RMSE | MAPE | |
NYCTaxi Pick-up CHBike Drop-off | 3.57 | 5.72 | 59.26% | 3.25 | 5.34 | 59.35% | 3.49 | 5.64 | 59.06% | 3.23 | 5.74 | 72.14% | 3.12 | 5.01 | 55.12% |
NYCTaxi Pick-up CHBike Pick-up | 3.61 | 5.98 | 59.55% | 3.29 | 5.60 | 59.71% | 3.53 | 5.91 | 59.14% | 3.25 | 5.15 | 88.52% | 3.06 | 5.40 | 50.94% |
NYCTaxi Pick-up NYCTaxi Drop-off | 9.99 | 20.22 | 75.14% | 10.00 | 21.16 | 68.03% | 11.03 | 21.86 | 70.32% | 11.02 | 22.34 | 94.31% | 9.31 | 18.68 | 66.42% |
NYCTaxi Drop-off CHBike Drop-off | 3.58 | 5.72 | 59.33% | 3.19 | 4.99 | 76.75% | 3.35 | 5.19 | 69.36% | 3.29 | 4.99 | 80.87% | 3.09 | 4.65 | 52.73% |
NYCTaxi Drop-off CHBike Pick-up | 3.62 | 5.99 | 59.55% | 3.26 | 5.27 | 79.41% | 3.43 | 5.49 | 71.76% | 3.33 | 5.32 | 82.60% | 3.02 | 5.18 | 68.27% |
NYCTaxi Drop-off NYCTaxi Pick-up | 10.04 | 17.72 | 88.10% | 9.67 | 17.76 | 73.46% | 8.09 | 14.58 | 50.99% | 11.14 | 20.57 | 94.03% | 8.02 | 13.21 | 46.16% |
5.6 Ablation Studies
With or Without Different Components. The ST-LLM’s effectiveness in traffic prediction stems from key components, explored through variant comparisons: w/o LLM (LLM removed), w/o ST (spatial-temporal embedding removed), w/o T (temporal embedding removed), and w/o S (spatial embedding removed). This analysis investigates the impact of each component. Figure 3 presents the ablation study on the NYCTaxi datasets. The w/o LLM variant shows a considerable increase in error across all metrics. Its removal leads to a degradation in performance, demonstrating that the prediction capabilities of the ST-LLM are heavily reliant on the LLM’s ability to learn complex dependencies from traffic data. The exclusion of ST results in a notable performance drop. This highlights the importance of spatial-temporal embedding in understanding both spatial and temporal dependencies in traffic data. The experimental results reveal that removing either temporal (w/o T) or spatial (w/o S) components similarly affects the model’s prediction error. Ablating either one of these embeddings results in an increased error. Notably, the model incurs a larger prediction error without the temporal component, underscoring the significance of our thoughtfully designed temporal embeddings.
Ablation Study of Partially Frozen Attention LLM. In this subsection, we conducted an ablation study to evaluate the efficacy of our proposed Partially Frozen Attention (PFA) LLM, as presented in Table 2. The PFA is compared against several variations: Frozen Pretrained Transformer (FPT), models without pretraining (No Pretrain), models utilizing the full 12 layers of GPT-2 (Full Layer), and fully tuned models without any frozen layers (Full Tuning). The PFA LLM demonstrates superior performance across all metrics on all datasets, which suggests that partially freezing the attention significantly enhances the predictive accuracy. FPT shows commendable performance, e.g., it is slightly outperformed by PFA. This indicates that the partial freezing strategy strikes a better balance between leveraging prelearned knowledge and adapting to new data. The Full Layer and Full Tuning models exhibit competitive performance. However, they still fall short of the accuracy demonstrated by the PFA model. The comparison with the No Pretrain model highlights the role of pretraining in model performance. While the No Pretrain model performs reasonably well, it is evident that pretraining, especially when combined with strategies like partial frozen is beneficial to achieving higher accuracy.
5.7 Parameter Analysis
Within the ST-LLM framework shown in Figure 4, represents the hyperparameter that determines the number of multi-head attention layers to unfreeze during the training process. As depicted in Figure 4 (a), an analysis of the NYCTaxi Pick-up dataset reveals that with an increase in , the WAPE initially decreases. The trend suggests that unfreezing additional layers up to a certain threshold can enhance the performance of the ST-LLM. However, this positive effect inverts when exceeds 2, at which point the model’s performance starts to degrade, hinting at the diminishing benefits associated with unfreezing more layers. In a nutshell, the optimal of taxi flow prediction is set to 2. Figure 4 (b) illustrates a similar pattern on the CHBike Pick-up dataset. The MAE indicates an increase in performance as is set to 1, with the lowest MAE observed at this value. This reinforces the observation that a single unfrozen multi-head attention layer is optimal for minimizing absolute errors, and the model achieves the balance between complexity and performance. This optimal point suggests that unfreezing more layers does not contribute to improved accuracy and might even degrade the performance of ST-LLM.
5.8 Few-Shot Prediction
In the few-shot prediction experiment, models trained with just 10% of data showed the ST-LLM model’s superior ability to recognize complex patterns from limited data, as detailed in Table 3. While the LLAMA2 model presents competitive results, especially on the NYCTaxi datasets, it does not consistently surpass the performance of ST-LLM. For instance, on the NYCTaxi pick-up dataset, ST-LLM achieves a noteworthy 7.06% reduction in MAE compared to LLAMA2. Other models, such as OFA, GATGPT, and GCNGPT, although commendable in their performances, do not match the superior results of ST-LLM. Notably, despite OFA’s better performance on the CHBike Drop-off dataset, ST-LLM still outperforms it with a 9.15% improvement in MAE. Compared with GATGPT and GCNGPT, ST-LLM shows remarkable average improvements of over 39.21% and 7.80% in MAE across all datasets, respectively. This significant difference highlights the robustness of ST-LLM in efficiently handling scenarios with limited data.
5.9 Zero-Shot Prediction
This zero-shot prediction is designed to assess the cross-dataset adaptation capability of LLMs. This task evaluates the model’s ability to generalize to the CHBike dataset after being trained exclusively on the NYCtaxi dataset, without any prior exposure to data from the CHBike dataset. The results of this evaluation are shown in Table 4. In terms of intra-domain transfer (e.g., from NYCTaxi Pick-up to NYCTaxi Drop-off), ST-LLM demonstrates its ability to maintain high accuracy, underlining its effectiveness in handling dataset-specific dependencies. The results also show that ST-LLM exhibits exceptional performance in inter-domain scenarios, such as transferring from NYCTaxi to CHBike. This is evident in all metrics, where ST-LLM consistently achieves the lowest error rates, indicating a robust ability to adapt to new domains without retraining. When analyzing other LLMs like OFA, GATGPT, GCNGPT, and LLAMA2, we observe varied performances. For instance, OFA performs well in certain scenarios but is generally outperformed by ST-LLM. GATGPT and GCNGPT, while showing competent adaptability, still fall short compared to ST-LLM’s performance, particularly in challenging inter-domain transfers. In conclusion, the zero-shot prediction results reinforce the adaptability and predictive strength of ST-LLM, making it a highly promising model for traffic flow prediction tasks across diverse datasets.
6 Conclusion
ST-LLM shows promise in adapting large language models for traffic prediction by embedding traffic data into spatial-temporal representations for LLMs. A partially frozen attention strategy within ST-LLM is proposed to enhance the model in traffic prediction. Our research findings demonstrate that the ST-LLM outperforms state-of-the-art models in traffic prediction. Looking forward, future work will explore a range of additional tasks, such as traffic imputation, and will focus on developing memory-efficient LLMs tailored for traffic data analysis.